Find all solutions using exact values cos 3.0 + cos 5x = 0 sec 2.c sec 63 = 0.
Answer #1

ricoset CosD so • Here given equation is cos3x + cos5x=0 ► 2005 ( 3x45x) cos (54,3%) = = 2008(et cos (CD) ?cos (2) ces (2) »> asin (8x) sin (4) = 0 - 2 sin(4x) sin (20) = 0 sin(4x) sin (2x) = 0 [:270] = sin(4x)=0 sin (2x)=0 4x = = nn 72x =nt x= 7 =